# The supermassive black hole population from seeding via collisions in Nuclear Star Clusters

### v1

@@ Appendix A Mass resolution and convergence tests
We checked the convergence of the Galacticus output running the same models listed in Table  1 changing the mass resolution.
@@ Appendix B Galaxy and black hole mass function deviations
Table 4: Summary of deviations for various models across different black hole mass ranges. (1) Model name; (2) average value of q of the model; (3) maximum value of q; (4) minimum value of q. All the quantities are in dex units.
@@ Appendix D Observed properties of galaxies hosting NSCs and SMBHs
Table 5: Properties of galaxies hosting NSCs and SMBHs. (1) Galaxy name; (2) black hole mass; (3) stellar mass of the NSC; (4) stellar mass of the host galaxy; (5) velocity dispersion of the bulge/pseudo-bulge/nuclei when corresponding; (6) references. All the columns show the estimated error when available in literature.
@@
(1) Genzel et al. (2010); (2) Schödel et al. (2014); (3) Bland-Hawthorn & Gerhard (2016); (4) Ferrarese & Merritt (2000); (5) Bender et al. (2005); (6) Kormendy & Ho (2013); (7) Williams et al. (2017); (8) Nguyen et al. (2019); (9) Nguyen et al. (2018); (10) Simien & Prugniel (2002); (11) van den Bosch & de Zeeuw (2010); (12) Nguyen et al. (2017); (13) Seth et al. (2010); (14) Barth et al. (2002); (15) Peterson & Caldwell (1993); (16) den Brok et al. (2015); (17) Reines & Volonteri (2015); (18) Filippenko & Ho (2003); (19) Krajnović et al. (2018); (20) Côté et al. (2006); (21) Cappellari et al. (2013); (22) Bower et al. (2001); (23) Graham & Spitler (2009); (24) Emsellem et al. (1999); (25) Gebhardt et al. (2003); (26) Schulze & Gebhardt (2011); (27) Gültekin et al. (2009a); (28) Saglia et al. (2016); (29) Gültekin et al. (2009b); (30) Lauer et al. (2005); (31) Cappellari et al. (2002); (32) Faber & Jackson (1976); (33) Pechetti et al. (2017); (34) Spengler et al. (2017); (35) Spolaor et al. (2010); (36) Gebhardt et al. (2001); (37) McConnachie (2012); (38) Quirk et al. (2022); (39) Neumayer & Walcher (2012); (40) Walcher et al. (2005); (41) Geha et al. (2002); (42) Barth et al. (2009); (43) Jarrett et al. (2003); (44) Satyapal et al. (2007); (45) Böker et al. (1999); (46) Karachentsev et al. (2013); (47) Graham & Soria (2019); (48) Gebhardt et al. (2011); (49) Bower et al. (1998); (50) Rusli et al. (2011); (51) Devereux et al. (2003); (52) Querejeta et al. (2015); (53) Williams (1977); (54) Ferrarese et al. (1996); (55) Shen & Gebhardt (2010); (56) de Francesco et al. (2006); (57) Nelson & Whittle (1995); (58) Sarzi et al. (2001); (59) Shapiro et al. (2006); (60) Burbidge et al. (1961); (61) Cretton & van den Bosch (1999).

### v2

@@ Appendix A Mass resolution and convergence tests
We checked the convergence of the Galacticus output running the same models listed in Table  4 changing the mass resolution.
@@ Appendix B Galaxy and black hole mass function deviations
Table 4: Summary of deviations for various models across different black hole mass ranges.
@@
121212(1) Model name; (2) average value of q of the model; (3) maximum value of q; (4) minimum value of q. All the quantities are in dex units.
@@ Appendix D Observed properties of galaxies hosting NSCs and SMBHs
Table 5: Properties of galaxies hosting NSCs and SMBHs.
@@
(1) Genzel et al. (2010); (2) Schödel et al. (2014); (3) Bland-Hawthorn & Gerhard (2016); (4) Ferrarese & Merritt (2000); (5) Bender et al. (2005); (6) Kormendy & Ho (2013); (7) Williams et al. (2017); (8) Nguyen et al. (2019); (9) Nguyen et al. (2018); (10) Simien & Prugniel (2002); (11) van den Bosch & de Zeeuw (2010); (12) Nguyen et al. (2017); (13) Seth et al. (2010); (14) Barth et al. (2002); (15) Peterson & Caldwell (1993); (16) den Brok et al. (2015); (17) Reines & Volonteri (2015); (18) Filippenko & Ho (2003); (19) Krajnović et al. (2018); (20) Côté et al. (2006); (21) Cappellari et al. (2013); (22) Bower et al. (2001); (23) Graham & Spitler (2009); (24) Emsellem et al. (1999); (25) Gebhardt et al. (2003); (26) Schulze & Gebhardt (2011); (27) Gültekin et al. (2009a); (28) Saglia et al. (2016); (29) Gültekin et al. (2009b); (30) Lauer et al. (2005); (31) Cappellari et al. (2002); (32) Faber & Jackson (1976); (33) Pechetti et al. (2017); (34) Spengler et al. (2017); (35) Spolaor et al. (2010); (36) Gebhardt et al. (2001); (37) McConnachie (2012); (38) Quirk et al. (2022); (39) Neumayer & Walcher (2012); (40) Walcher et al. (2005); (41) Geha et al. (2002); (42) Barth et al. (2009); (43) Jarrett et al. (2003); (44) Satyapal et al. (2007); (45) Böker et al. (1999); (46) Karachentsev et al. (2013); (47) Graham & Soria (2019); (48) Gebhardt et al. (2011); (49) Bower et al. (1998); (50) Rusli et al. (2011); (51) Devereux et al. (2003); (52) Querejeta et al. (2015); (53) Williams (1977); (54) Ferrarese et al. (1996); (55) Shen & Gebhardt (2010); (56) de Francesco et al. (2006); (57) Nelson & Whittle (1995); (58) Sarzi et al. (2001); (59) Shapiro et al. (2006); (60) Burbidge et al. (1961); (61) Cretton & van den Bosch (1999). 131313(1) Galaxy name; (2) black hole mass; (3) stellar mass of the NSC; (4) stellar mass of the host galaxy; (5) velocity dispersion of the bulge/pseudo-bulge/nuclei when corresponding; (6) references. All the columns show the estimated error when available in literature.